The Fashion Shoe Company operates a chain of womens shoe shops that carry many styles of shoes that are all sold at the same price. Sales personnel in the shops are paid a sales commission on each pair of shoes sold plus a small base salary. The following data pertains to Shop 48 and is typical of the companys many outlets Per Pair of Shoes $ 30.00 Selling price Variable expenses: $ 13.50 Invoice cost Sales commission 4.50 $ 18.0e Annual Total variable expenses Fixed expenses: Advertising Rent Salaries $30,000 20,000 100,000 $150,0060 Total fixed expenses

5. Refer to the original data. As an alternative to (4) above, the company is considering paying the Shop 48 store manager 50 cents commission on each pair of shoes sold in excess of the break-even point. If this change is made, what will be Shop 48s net operating income (loss) if 15,000 pairs of shoes are sold? (Do not round intermediate calculations.) Answer is complete but not entirely correct. Net operatings 22,500 income

6. Refer to the original data. The company is considering eliminating sales commissions entirely in its shops and increasing fixed salaries by $31,500 annually. If this change is made, what will Shop 48s new break-even point in unit sales and dollar sales? (Do not round intermediate calculations.) Answer is complete but not entirely correct. New break-even point in unit sales 315,000pairs New break-even point in dollar sales $ 572,700

Need help with the wrong answers
Answer #1
Part 5)
The revised net operating income/loss can be calculated with the use of following formula:
Revised Net Operating Income/Loss = Sales in Excess of Break Even*(Selling Price - Revised Variable Cost) Break even sales= 150000/ (30-18) 12500
Revised Variable Cost = 18 +0.50 = $18.50
Revised Net Operating Income/Loss = (15,000 - 12,500)*(30 - 18.50) = $28,750
____________
Part 6a)
Revised Fixed Cost = 30,000 + 20,000 + (100,000 + 31500) = $181,500
Revised Variable Cost = $ 13.50 per unit [commission completely eliminated]
____________
Using these values in the formulas for break even point, we get,
Revised Break Even Point (Units) = (181500)/(30 - 13.50) = 11,000 pairs
Revised Break Even Point (Dollars) = 11000*30 = $330,000